OpenStudy (anonymous):

Find the distance between the points (-1,4) and (-2,-9).

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

hartnn (hartnn):

no?

OpenStudy (anonymous):

yes

hartnn (hartnn):

y do i feel that u are writing for last 10 minutes...

OpenStudy (anonymous):

ohh sorry i took a shoert

OpenStudy (anonymous):

shower

hartnn (hartnn):

great! so could u find the distance ?

OpenStudy (anonymous):

idk

hartnn (hartnn):

what is -1-(-2) = ?

OpenStudy (anonymous):

-3

hartnn (hartnn):

nopes

OpenStudy (anonymous):

-1

hartnn (hartnn):

-1-(-2) = -1+2 = ?

OpenStudy (anonymous):

1

hartnn (hartnn):

4-(-9) = ?

OpenStudy (anonymous):

5

hartnn (hartnn):

4-(-9) = 4+ 9 = 13

hartnn (hartnn):

now what is 13^2 + 1^2 = ?

hartnn (hartnn):

then take its square root

OpenStudy (anonymous):

169+2

hartnn (hartnn):

169+1 = 170 answer = \(\sqrt{170}\)

OpenStudy (anonymous):

ohh ok i get it but i don't get like 4-(-9)

hartnn (hartnn):

- * - = + so 4-(-9) = 4+ 9 = 13